OpenStudy (anonymous):

Please help!! I will give you a medal if you help me. Given the equation y - 4 = 3/4(x + 8) in point-slope form, identify the equation of the same line in standard form. -3/4x + y = 10 3x - 4y = -40 y = 3/4x + 12 y = 3/4x + 10

OpenStudy (texaschic101):

y - 4 = 3/4(x + 8) -- distribute through the parenthesis y - 4 = 3/4x + 6 -- add 4 to both sides y = 3/4x + 6 + 4 y = 3/4x + 10 now we turn to standard form : Ax + By = C y = 3/4x + 10 -- multiply the equation by the LCD of 4 4y = 3x + 40 -- subtract 3x from both sides -3x + 4y = 40 -- multiply by -1 to make x positive 3x - 4y = -40 <== standard form

OpenStudy (anonymous):

Oh my gosh!! Thank you sooo much!!!

OpenStudy (texaschic101):

no problem :)